find the sum of each series
-60 + (-59) + (-58) + . . . . + 87 + 88 + 89 + 90

Answer by stanbon(75887)

find the sum of each series
-60 + (-59) + (-58) + . . . . + 87 + 88 + 89 + 90
----
Formula: S(n) = (n/2)(1st + last)
n = 90+1+60 = 151 terms
-------
S(151) = (151/2)(-60+90)
= 75.5*30
= 2265
================
Cheers,
Stan H.